OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of the function f(x)=(x−2)(x−7)3+7 on each of the indicated intervals. (a) Interval = [1,4]. 1. Absolute maximum = 2. Absolute minimum = (b) Interval = [1,8]. 1. Absolute maximum = 2. Absolute minimum = (c) Interval = [4,9]. 1. Absolute maximum = 2. Absolute minimum =

myininaya (myininaya):

Hey.

myininaya (myininaya):

You can find the derivative of this right?

OpenStudy (anonymous):

yeah

myininaya (myininaya):

Is that \[f(x)=(x-2)(x-7)^3+7\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the derivative is (x-7)^2 (4 x-13)

myininaya (myininaya):

we are finding the derivative to find the critical numbers you will need to solve f'(x)=0 for x we don't have to find when f' dne since this is a polynomial and polynomials are differentiable everywhere

myininaya (myininaya):

i'm going to check your derivative before we move on

myininaya (myininaya):

looks fabulous

myininaya (myininaya):

ok now that is very first step in problem like i will get you the recipe 1) find derivative 2) check if f is continuous on the domain given (don't worry here; polys are continuous everywhere) 3) find for what x is the derivative zero solve f'=0 for x 4) find for what x is the derivative not existing (don't worry here; polys are differentiable everywhere) 5) then evaluate f(endpoints) and f(critical numbers) 6) determine which values are highest and which are lowest in 5 for answer

myininaya (myininaya):

so we need to find when (x-7)^2(4x-13)=0

myininaya (myininaya):

when is x-7=0? when is 4x-13=0?

myininaya (myininaya):

oh also on number 5 you only want to plug in critical numbers that are between the endpoints

OpenStudy (anonymous):

so whats the first thing after the derivative?

myininaya (myininaya):

we are trying to find when the derivative is zero

myininaya (myininaya):

that is what i set the derivative equal to zero and asked you to solve it

myininaya (myininaya):

do you know how to solve (x-7)^2(4x-13)=0?

myininaya (myininaya):

recall if you have a*b=0 this implies a=0 or b=0 or both =0

myininaya (myininaya):

so x-7=0 or 4x-13=0 so x=? or x=?

OpenStudy (anonymous):

i just did that and theres no number that would make 4x-13 equal to 0, and the answer for the other one is 7

myininaya (myininaya):

you can solve 4x-13=0

myininaya (myininaya):

come on you are taking calculus you have to recall some algebra for it

myininaya (myininaya):

if you want to get the x by itself you need to do two operations to both sides

myininaya (myininaya):

and i say that because there is subtraction to undo and also multiplication

OpenStudy (anonymous):

ok, the number is 3.25

myininaya (myininaya):

4x-13=0 +13 +13 <--add 13 to both sides 4x =13 -- -- <---divide 4 on both sides 4 4 x=13/4 you also found the other critical number x=7 we have two critical numbers in all we will found out which ones matter when we look at the intervals

myininaya (myininaya):

So the first function is defined [1,4] we only care about the function from x=1 to x=4

myininaya (myininaya):

but guess what x=7 doesn't occur between there so we don't care about x=7 for the first part.

OpenStudy (anonymous):

ohh but the second one does right?

myininaya (myininaya):

now go to number 5 of my recipe evaluate f(endpoints) and f(critical numbers) make sure to only plug in critical number we care about for this one

myininaya (myininaya):

yes and we will get to that one

myininaya (myininaya):

you can go ahead and evaluate f(7) and f(13/4) but we will not use f(7) for the first one

myininaya (myininaya):

so what is f(1) and f(4) and f(13/4)

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

(x-7)^2 (4 x-13) when we use 1, we get -324 when we use 4, we get 27 when we use 13/4, we get 0

myininaya (myininaya):

did you find f(1) and f(4) and f(13/4)?

myininaya (myininaya):

i need those values

OpenStudy (anonymous):

yeah

myininaya (myininaya):

i don't care where the derivative it at that number

myininaya (myininaya):

i care about where they occur on the graph i'm trying to determine the max and the min

myininaya (myininaya):

ok can you give me those values...

myininaya (myininaya):

Recall f(x)=(x-2)(x-7)^3+7 to find f(1) you replace the x's with 1 like so f(1)=(1-2)(1-7)^3+7

myininaya (myininaya):

then evaluate and simplify

OpenStudy (anonymous):

ohh my bad i took them in the derivative

OpenStudy (anonymous):

(x−2)(x−7)^3+7 at 1, it is 223 at 4, its -47 at 13/4, -58.91

myininaya (myininaya):

groovy now which of those is the highest value (absolute max) which of those is the smallest value (absolute min)

OpenStudy (anonymous):

223 absolute max and -58.91 absolute min

myininaya (myininaya):

and i like exact numbers so instead of -58.92 i would say -15083/256

myininaya (myininaya):

ok anyways you should be able to do the other problems they are the same exact way

OpenStudy (anonymous):

thank you

myininaya (myininaya):

go ahead and do the 2nd one and let me know what you get

myininaya (myininaya):

also the 2nd one is tons easier because we already did most of the work

myininaya (myininaya):

you don't need to find f' again you don't need to find the critical numbers x=13/4 or x=7

myininaya (myininaya):

you also don't need to evaluate f(1) again

myininaya (myininaya):

or f(13/4)

OpenStudy (anonymous):

ok will do

myininaya (myininaya):

now i said you don't need to find them but you still need to use the info

OpenStudy (anonymous):

i found the other two 2) max: 223 and min: -58.91 3) max: 63 and min: -47

myininaya (myininaya):

looks good

myininaya (myininaya):

great job proto